# Emergence of Hidden Capabilities: Exploring Learning Dynamics in Concept Space

### v1

@@ 1 Introduction
Figure 1: Concept Learning Geometry underlies emergence. (a) Top: A multimodal model learns to generate the concepts in the order of “astronaut”, “horse”, and finally “riding” as it scales up (adapted from Yu et al.[45]). Middle: “blue square apple” is generated in the order of “apple”, “square”, and “blue” (adapted from Li et al.[46]). Bottom: Despite its simplicity, our model trained on synthetic data shows concept learning dynamics where it first learns “shape” and then “color”. (b) Concept space is an abstract coordinate space where individual axes correspond to different concepts and a given point corresponds to a “concept class”, i.e., a predefined collection of concepts (e.g., large blue circles on bottom left corner). Traversal along axes of the concept space yield change in a specific property of the sample (e.g., going from large blue circle to large red circle along object color axis). Trajectories show a model’s dynamics in concept space for learning to generate classes shown in-distribution (blue nodes) versus out of distribution (pink nodes). As we show, dynamics in concept space are highly interpretable, enabling precise comments on which concepts the model learns first, why, and what order it follows. (c) Measuring how accurately a model generates samples from a given concept class, showing an order of concept learning: first background color is learned, then size, and then object color.
@@ 2 Related Work
Concept learning. The term concepts as used in this work is broadly equivalent to the notion of factors of variation from prior work on disentangled representation learning [23, 54, 55, 56, 57, 30, 58, 59, 60], and is motivated by use of the term in a similar sense in cognitive science [61, 62, 63, 64, 65, 66]. The focus of literature on disentanglement has been to prove identifiability results, e.g., when will a generative model trained on a dataset learn to invert the data-generating process, hence identifying the latent variables, i.e., concepts, that underlie it. Given the success of modern generative models, we argue, despite impossibility results from prior work, these models are in fact learning to disentangle some, if not all, concepts. Precisely what guides which concepts are disentangled however requires studying the learning dynamics—the target of our work.
Interpretability. A growing line of papers demonstrate highly interpretable representations of intuitive concepts exist in generative models, especially LLMs [67, 68, 69, 70, 71, 72]. Similar work in this vein in image diffusion models has found semantic representations in various components of the model [52, 73, 74, 75, 37]: e.g., existence of linear representations for concepts such as 3D depth or object versus background distinctions [76, 77] . These papers further bolster our argument, modern generative models are truly inverting the data-generating process and identifying the concepts underlying it.
@@ 3 Concept Space: A Framework for Analyzing Concept Learning Dynamics
We first formally introduce our framework of concept spaces. Motivated by a rich body of literature on disentangled representation learning [23, 54, 55, 56, 57, 30, 58, 59, 60], this framework allows us to systematically analyze how different concepts underlying the data generating process and corresponding capabilities to manipulate them are learned during training. We highlight that since our primary empirical focus in this paper will be an abstraction of text-to-image generative models, parts of the framework will be specifically instantiated with text-to-image generation tasks in mind. To this end, we note our model class of interest is a generative model $F$ that is trained using conditioning information $h$ to produce images $y$. For example, $F$ may be instantiated using a diffusion model that uses embeddings $h$ of textual descriptions of a scene to produce images $y$. We next define a concept space.
@@ 3.1 Experimental and Evaluation Setup
To assess whether a generated image matches the desired concept class without human intervention, we follow literature on disentangled representation learning [89, 54, 90, 91, 92, 55, 23, 30] and train classifier probes for individual concepts using the diffusion model’s training data. The probe architecture involves a U-Net [93] followed by an average pooling layer and $n$ MLP classification heads for the $n$ concept variables. See App. A.2 for further details.
@@ 4.1 Concept Signal Determines Learning Speed
We first demonstrate the utility of concept signal as a tool to gauge at what rate the model learns a concept and the capability to manipulate it. To this end, we develop controlled variants of our data-generating process by changing the level of concept signal of each concept and train diffusion models conditioned with the latent concept vector $z$ on them. We primarily focus on concepts color and size, altering their concept signal by respectively adjusting the RGB contrast between red and blue and the size difference between large and small objects (see App. A.1 for details). We define the speed of learning each concept as inverse of the number of gradient steps required to reach 80% accuracy for class 11, i.e., the OOD class that requires learning the capability to manipulate concepts as seen during training. Results comparing different concept signals are shown in Fig. 3. For both color and size, we observe that concept signal dictates the speed at which individual concepts are learned. We also find that when different concepts have varying strengths of concept signals, this leads to differences in the learning speed for each concept.
Figure 4: Concept signal governs generalization dynamics. (a) Learning dynamics in the concept space for in-distribution concept class 00 (bottom left). (b) Learning dynamics for out-of-distribution (OOD) concept class 11 (top right). We plot the accuracy for color on the x-axis and size on the y-axis. The [0,1) normalized color concept signal level is color coded. Two trajectories for 01 and 10 are shown to illustrate concept memorization.
@@ 4.2 Concept Signal Governs Generalization Dynamics
We next examine the model’s learning dynamics in concept space under various levels of concept signal for the concepts color and size. For completeness, we evaluate a model’s ability to generalize both in-distribution (ID) and OOD, but only the latter is deemed learning of a capability to manipulate a concept. Results are shown in Fig. 4, with panel (a) showing dynamics for learning to generate samples from ID class 00 and panel (b) showing dynamics for OOD class 11. Results on OOD generalization show a form of concept memorization, which we define as the phenomenon where the model’s generations on an OOD conditioning $h$ are biased towards the training class that helps define the strongest concept signal. For example, for the unseen conditioning 11, the generations are more alike 01 when the concept signal is stronger for size (e.g., blue curve in Fig. 4(b)) and alike 10 when the signal is stronger for color (e.g., red curve). Interestingly, we observe that for settings with high imbalance in concept signals, e.g., the blue curve in Fig. 4 (b), the endpoint of concept memorization is very biased towards one training class, here 01, delaying its out-of-distribution (OOD) generalization. For the learning trajectories of all classes, see Fig. 13 in App. C.2.
Broadly, our results imply that an early stopped text-to-image model can witness concept memorization and hence simply associate an unseen conditioning to the nearest concept class when asked to generate OOD samples (see Kang et al. [94] for a similar result in LLMs). However, given sufficient time, the model will disentangle concepts underlying the data-generating process and learn to generate entirely novel, OOD samples. For futher evidence in this vein, we also confirm our results across more general scenarios, including with the real-world CelebA dataset (App. C.3) and using three concept variables: color, background color, and size (App. C.4).
@@ 4.4 Sudden Transitions in Concept Learning Dynamics
The concept space visualization of learning dynamics observed in Fig. 4 (b) and our toy models analysis in Fig. 5 indicate that there exists a phase in which the model departs from concept memorization and disentangles each of the concepts, but still produces incorrect images. We claim that at the point of departure, the model has in fact already disentangled concepts underlying the data-generating process and acquired the relevant capabilities to manipulate them, hence yielding a change of direction in its learning trajectory. However, naive input prompting is insufficient to elicit these capabilities and generate samples from OOD classes, giving the impression the model is not yet “competent”. This then leads to the second phase in the learning dynamics, wherein an alignment between the input space and underlying concept representations is learned. We take the model corresponding to the second from left curve (the green curve) in Fig. 4 (b) to investigate this claim in detail. Specifically, given intermittent checkpoints along the model’s learning trajectory, we use the following two protocols for prompting the model to produce images corresponding to the class 11 (blue, small). See App. B for further details.
@@
    Input Space: Overprompting. We simply enhance the color conditioning to values of higher magnitude, e.g. (0.4, 0.4, 0.6) to (0.3, 0.3, 0.7).
@@
We further explore the second phase of learning in Appendix C.6 by patching the embedding module used for processing the conditioning information from the final checkpoint to an intermediate one. Our results show that when the final checkpoint does enable use of naive input prompting for eliciting a capability, the embedding module can be patched to an intermediate checkpoint and we can retrieve the desired output at approximately the same time that alternative prompting protocols start to work well. This suggests the second phase of learning primarily involves aligning the input space to intermediate representations that enable eliciting the model capabilities. Overall, our results above yield the following hypothesis.
@@ 5 Effect of Underspecification on Learning Dynamics in Concept Space
Experimental setup. The data generation and evaluation process follows the protocol described in Sec. 3.1. We select color (red and blue) and shape (circle and triangle) as the concepts, drawing an analogy to the “yellow strawberry” example. To simulate underspecification, we randomly select training samples that have a specific combination of shape and color (e.g., “red triangle”). We then mask the token representing the color (“red”) and train the model on three concept classes {00, 01, 10}, represented by blue nodes in Fig. 7, with some prompts masked. We test using class 11 (pink node) with no prompts masked, to see if the model has learned disentangled concepts and can generalize OOD.
@@
$\frac{d{\bf z}}{dt}=-\nabla_{\bf z}L,\,\,L(z_{1},z_{2})=\big{(}(1-s\alpha)-z_{1}\big{)}^{2}+\big{(}1-z_{2}\big{)}^{2}.$ (2)
@@
Influence of Underspecification on Emergence of Hidden Capabilities. Following Sec. 4.4, we also explore whether it is possible to elicit the desired outputs from a model trained with underspecified data. Fig. 10 shows the accuracy results over five runs (a) without using any prompting method, and (b) using over-prompting. In both scenarios, the percentage of masking is set at 50%. Results clearly demonstrate that with over-prompting, the model achieves 100% accuracy after approximately 1,000 gradient steps, whereas without over-prompting, it fails in three out of five runs even after 2,000 gradient steps. These findings confirm that capability can develop prior to observable behavior, even in cases of underspecification.
@@ 6 Discussion
Why study the concept space? One might ask the question of why the concept space could be useful beyond synthetic datasets. In Fig. 12, we show the loss, accuracy, and training trajectory in concept space of a model from Fig. 6. The time at which the model acquires the capability to manipulate size and color independently is not evident from either the loss or accuracy curves. However, in concept space, one can directly see the divergence of the trajectory from concept memorization. Benchmarking generative models is a challenging task, still often involving humans in the loop [102, 103]. Our concept space framework suggests that benchmarking out-of-distribution (OOD) generalization can potentially be reduced to monitoring the learning trajectory in concept space.
Concept Learning vs. Grokking. We make a distinction between our delayed elicitation of capabilities versus grokking [104]. The common aspect is that the measured performance on a test set is delayed from the train set, as seen in Fig. 12. However, we deal with out-of-distribution test data, which is different from the setups like modular arithmetic or polynomial regression in which grokking is usually explored [104, 105, 106, 107, 108]. Second, work on hidden progress measures show that in fact a model is continuously building representations towards the mechanism to solve the specified task [107, 109]; however, our results find that even at the level of latent representations, there is a sudden emergence whereby the model learns the capability to manipulate concepts underlying the data-generating process and generalize OOD.
@@
In Sec. 4.4, we have demonstrated that before the capability to manipulate a concept is learned, the desired outputs cannot be generated irrelevant of the protocol used for prompting the model. Moreover, we showed that different protocols yield desired outputs at precisely the same time, and this time is also independent of model initialization. We thus hypothesize that: Concept learning is a well-controlled phase transition in model capability, but the ability to elicit this capability via a predefined single input prompting can be arbitrarily delayed, e.g., by the level of the concept signal.
@@
CFP and HT gratefully acknowledges the support of Aravinthan D.T. Samuel. CFP thanks Zechen Zhang for useful discussions. ESL’s time at University of Michigan was partially supported by the National Science Foundation (IIS-2008151) and at CBS, Harvard by NTT Research, Inc. The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
Appendix A Experimental Details
A.1 Synthetic data

### v2

@@ 1 Introduction
Figure 1: Concept Learning Geometry underlies emergence. (a) Top: A multimodal model learns to generate the concepts in the order of “astronaut”, “horse”, and finally “riding” as it scales up (adapted from Yu et al.[45]). Middle: “blue square apple” is generated in the order of “apple”, "blue", and “square”” (adapted from Li et al.[46]). Bottom: Despite its simplicity, our model trained on synthetic data shows concept learning dynamics where it first learns “shape” and then “color”. (b) Concept space is an abstract coordinate space where individual axes correspond to different concepts and a given point corresponds to a “concept class”, i.e., a predefined collection of concepts (e.g., large blue circles on bottom left corner). Traversal along axes of the concept space yield change in a specific property of the sample (e.g., going from large blue circle to large red circle along object color axis). Trajectories show a model’s dynamics in concept space for learning to generate classes shown in-distribution (blue nodes) versus out of distribution (pink / red nodes). As we show, dynamics in concept space are highly interpretable, enabling precise comments on which concepts the model learns first, why, and what order it follows. (c) Measuring how accurately a model generates samples from a given concept class, showing an order of concept learning: first background color is learned, then size, and then object color.
@@ 2 Related Work
Concept learning. The term concepts as used in this work is broadly equivalent to the notion of factors of variation from prior work on disentangled representation learning [23, 54, 55, 56, 57, 30, 58, 59, 60, 61], and is motivated by use of the term in a similar sense in cognitive science [62, 63, 64, 65, 66, 67]. The focus of literature on disentanglement has been to prove identifiability results, e.g., when will a generative model trained on a dataset learn to invert the data-generating process, hence identifying the latent variables, i.e., concepts, that underlie it. Given the success of modern generative models, we argue, despite impossibility results from prior work, these models are in fact learning to disentangle some, if not all, concepts. Precisely what guides which concepts are disentangled however requires studying the learning dynamics—the target of our work.
Interpretability. A growing line of papers demonstrate highly interpretable representations of intuitive concepts exist in generative models, especially LLMs [68, 69, 70, 71, 72, 73, 74]. Similar work in this vein in image diffusion models has found semantic representations in various components of the model [52, 75, 76, 77, 37]: e.g., existence of linear representations for concepts such as 3D depth or object versus background distinctions [78, 79] . These papers further bolster our argument, modern generative models are truly inverting the data-generating process and identifying the concepts underlying it.
@@ 3 Concept Space: A Framework for Analyzing Concept Learning Dynamics
We first formally introduce our framework of concept spaces. See Fig. 12 for a schematic. Motivated by a rich body of literature on disentangled representation learning [23, 54, 55, 56, 57, 30, 58, 59, 60], this framework allows us to systematically analyze how different concepts underlying the data generating process and corresponding capabilities to manipulate them are learned during training. We highlight that since our primary empirical focus in this paper will be an abstraction of text-to-image generative models, parts of the framework will be specifically instantiated with text-to-image generation tasks in mind. To this end, we note our model class of interest is a generative model $F$ that is trained using conditioning information $h$ to produce images $y$. For example, $F$ may be instantiated using a diffusion model that uses embeddings $h$ of textual descriptions of a scene to produce images $y$. We next define a concept space.
@@ 3.1 Experimental and Evaluation Setup
To assess whether a generated image matches the desired concept class without human intervention, we follow literature on disentangled representation learning [91, 54, 92, 93, 94, 55, 23, 30] and train classifier probes for individual concepts using the diffusion model’s training data. The probe architecture involves a U-Net [95] followed by an average pooling layer and $n$ MLP classification heads for the $n$ concept variables. See App. B.2 for further details.
@@ 4.1 Concept Signal Determines Learning Speed
We first demonstrate the utility of concept signal as a tool to gauge at what rate the model learns a concept and the capability to manipulate it. To this end, we develop controlled variants of our data-generating process by changing the level of concept signal of each concept and train diffusion models conditioned with the latent concept vector $z$ on them. We primarily focus on concepts color and size, altering their concept signal by respectively adjusting the RGB contrast between red and blue and the size difference between large and small objects (see App. B.1 for details). We define the speed of learning each concept as inverse of the number of gradient steps required to reach 80% accuracy for class 11, i.e., the OOD class that requires learning the capability to manipulate concepts as seen during training. Results comparing different concept signals are shown in Fig. 3. For both color and size, we observe that concept signal dictates the speed at which individual concepts are learned. We also find that when different concepts have varying strengths of concept signals, this leads to differences in the learning speed for each concept.
Figure 4: Concept signal governs generalization dynamics. (a) Learning dynamics in the concept space for in-distribution concept class 00 (bottom left). (b) Learning dynamics for out-of-distribution (OOD) concept class 11 (top right). We plot the accuracy for color on the x-axis and size on the y-axis. The [0,1) normalized color concept signal level is color coded. Two trajectories for 01 and 10 are shown to illustrate concept memorization. See App. D.3 for uncertainties.
@@ 4.2 Concept Signal Governs Generalization Dynamics
We next examine the model’s learning dynamics in concept space under various levels of concept signal for the concepts color and size. For completeness, we evaluate a model’s ability to generalize both in-distribution (ID) and OOD, but only the latter is deemed learning of a capability to manipulate a concept. Results are shown in Fig. 4, with panel (a) showing dynamics for learning to generate samples from ID class 00 and panel (b) showing dynamics for OOD class 11. Results on OOD generalization show concept memorization, which we define as the phenomenon where the model’s generations on an OOD conditioning $h$ are biased towards the training class that helps define the strongest concept signal. For example, for the unseen conditioning 11, the generations are more alike 01 when the concept signal is stronger for size (e.g., blue curve in Fig. 4(b)) and alike 10 when the signal is stronger for color (e.g., red curve). Interestingly, we observe that for settings with high imbalance in concept signals, e.g., the blue curve in Fig. 4 (b), the endpoint of concept memorization is very biased towards one training class, here 01, delaying its out-of-distribution (OOD) generalization. For the learning trajectories of all classes, see Fig. 15 in App. D.2.
Broadly, our results imply that an early stopped text-to-image model can witness concept memorization and hence simply associate an unseen conditioning to the nearest concept class when asked to generate OOD samples (see Kang et al. [96] for a similar result in LLMs). However, given sufficient time, the model will disentangle concepts underlying the data-generating process and learn to generate entirely novel, OOD samples. For futher evidence in this vein, we also confirm our results across more general scenarios, including with the real-world CelebA dataset (App. D.4) and using three concept variables: color, background color, and size (App. D.5).
@@ 4.4 Sudden Transitions in Concept Learning Dynamics
The concept space visualization of learning dynamics observed in Fig. 4 (b) and our toy models analysis in Fig. 5 indicate that there exists a phase in which the model departs from concept memorization and disentangles each of the concepts, but still produces incorrect images. We claim that at the point of departure, the model has in fact already disentangled concepts underlying the data-generating process and acquired the relevant capabilities to manipulate them, hence yielding a change of direction in its learning trajectory. However, naive input prompting is insufficient to elicit these capabilities and generate samples from OOD classes, giving the impression the model is not yet “competent”. This then leads to the second phase in the learning dynamics, wherein an alignment between the input space and underlying concept representations is learned. We take the model corresponding to the second from left curve (the green curve) in Fig. 4 (b) to investigate this claim in detail. Specifically, given intermittent checkpoints along the model’s learning trajectory, we use the following two protocols for prompting the model to produce images corresponding to the class 11 (blue, small). See App. C for further details.
@@
    Input Space: Overprompting. We simply enhance the color conditioning to values of higher magnitude, e.g. $(r,\>g,\>b)$ = (0.4, 0.4, 0.6) to (0.3, 0.3, 0.7).
@@
We further explore the second phase of learning in Appendix D.7 by patching the embedding module used for processing the conditioning information from the final checkpoint to an intermediate one. Our results show that when the final checkpoint does enable use of naive input prompting for eliciting a capability, the embedding module can be patched to an intermediate checkpoint and we can retrieve the desired output at approximately the same time that alternative prompting protocols start to work well. This suggests the second phase of learning primarily involves aligning the input space to intermediate representations that enable eliciting the model capabilities. Overall, our results above yield the following hypothesis.
@@
4.5 Additional Results on Realistic Data
To provide further support for our hypothesis, we provide results in a more realistic and naturalistic data. Specifically, we use the CelebA dataset [98], which contains fine-grained attributes corresponding to concepts like Gender, With Hat, and Smiling, and analyze two settings.
  • Using concepts Gender and Smiling. Results are in Fig. 18, where we find that concept learning dynamics discovered in the relatively simple setting in prior sections generalizes. We see that there is first a phase of concept memorization, wherein the model learns to generate images of (FemaleSmiling). However, as training proceeds, there is a sudden turn in the learning dynamics and the model learns to generalize out-of-distribution.
  • Using concepts Gender and With Hat. In this setting, given a compute budget of 2M gradient steps, we find that the model seem to never learn to generalize out-of-distribution (see Fig. 7). However, when we perform latent interventions on the model’s representations, we are able to force the model to produce images from the class (Female, With Hat). These results reiterate that the model is indeed capable of generalizing out-of-distribution, but is not performant when naively prompted.
Figure 7: Validating our Findings on CelebA. (a) Concept space dynamics of generated images concepts Gender and With Hat. We find the generalization class (Female, With Hat) ends up near (Female, No Hat). (b) Compositional Generalization Accuracy when performing latent interventions vs. naive prompting. This rise of accuracy near $5\times 10^{5}$ gradient steps clearly demonstrates that the model is capable of generalizing out-of-distribution, but is not performant when evaluated via naive prompting.
@@ 5 Effect of Underspecification on Learning Dynamics in Concept Space
Experimental setup. The data generation and evaluation process follows the protocol described in Sec. 3.1. We select color (red and blue) and shape (circle and triangle) as the concepts, drawing an analogy to the “yellow strawberry” example. To simulate underspecification, we randomly select training samples that have a specific combination of shape and color (e.g., “red triangle”). We then mask the token representing the color (“red”) and train the model on three concept classes {00, 01, 10}, represented by blue nodes in Fig. 9, with some prompts masked. We test using class 11 (pink node) with no prompts masked to see if the model can generalize OOD.
@@
$\small\frac{d{\bf z}}{dt}=-\nabla_{\bf z}L,\,\,L(z_{1},z_{2})=\big{(}(1-s\alpha)-z_{1}\big{)}^{2}+\big{(}1-z_{2}\big{)}^{2}.$ (2)
@@
Influence of Underspecification on Emergence of Hidden Capabilities.
Following Sec. 4.4, we also explore whether it is possible to elicit the desired outputs from a model trained with underspecified data. Fig. 11 shows the accuracy results over five runs (a) without using any prompting method, and (b) using over-prompting. In both scenarios, the percentage of masking is set at 50%. Results clearly demonstrate that with over-prompting, the model achieves 100% accuracy after approximately 1,000 gradient steps, whereas without over-prompting, it fails in three out of five runs even after 2,000 gradient steps. These findings confirm that capability can develop prior to observable behavior, even in cases of underspecification.
@@ 6 Discussion
Why study the concept space? One might ask why concept space could be useful beyond synthetic datasets. In Fig. 14, we show the loss, accuracy, and training trajectory in concept space of a model from Fig. 6. The moment during training when the model acquires the capability to manipulate size and color independently is not evident from either the loss or accuracy curve. However, in concept space, one can directly see the divergence of the trajectory from concept memorization. Benchmarking generative models is a challenging task, still often involving humans in the loop [105, 106]. Our concept space framework suggests that benchmarking core generalization capabilities can potentially be reduced to monitoring the learning trajectory in concept space. Moreover, this information can be used to develop better training schemes as discussed in Ren et al. [107].
Concept learning vs. grokking
We make a distinction between the potentially delayed elicitation of capabilities in our model versus grokking [108]. The phenomenology of delayed increase in performance on the test set, as seen in Fig. 14, is shared. However, we deal with out-of-distribution evaluations, which is different from setups like modular arithmetic or polynomial regression in which grokking is usually studied [108, 109, 110, 111, 112]. Secondly, research on grokking using hidden progress measures indicates that the model gradually builds representations toward an ideal one [111, 113]; however, our results find that even at the level of latent representations, there is a sudden emergence whereby the model learns the capability to manipulate concepts underlying the data-generating process and generalize OOD.
@@
In Sec. 4.4, we have demonstrated that before the capability to manipulate a concept is learned, the desired outputs cannot be generated regardless of the protocol used to prompt the model. Moreover, we showed that different protocols yield the desired outputs at the exact same time, which is also highly independent of model initialization (Fig. 6, Fig. 17). We thus hypothesize that: Concept learning is a well-controlled phase transition in model capability. However, the ability to elicit this capability through predefined, single-input prompting can be delayed arbitrarily, depending on factors like the strength of the concept signal.
Limitations
One limitation of our work is that the compositional setup used in this study is rather simple. Real world concepts are often hierarchical [114, 115] or relational [116]. We also make assumptions that concepts are linearly embedded in the vector space $\mathcal{Z}$ in Sec. 3. However, this assumption seems to be at least partially justified as seen in Sec. D.8. Another limitation is that the findings are mainly drawn from synthetic data. While we show that our findings generalize to realistic data to some extent, as seen in Sec 4.5, Fig. 18 and Fig. 7, further studies are needed to investigate how our findings apply to real data in general.
@@
CFP and HT gratefully acknowledges the support of Aravinthan D.T. Samuel. CFP, MO, ESL and HT are supported by NTT Research under the CBS-NTT Physics of Intelligence program. ESL’s time at University of Michigan was partially supported by the National Science Foundation (IIS-2008151 and CNS-2211509). The computations in this paper were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University. The authors thanks Zechen Zhang, Helena Casademunt, Carolina Cuesta-Lazaro, Shivam Raval, Yongyi Yang and Itamar Pres for useful discussions.
Contributions
Motivated by prior work from MO, ESL and HT [30], CFP and HT started investigating the order in which concepts are learned first by a diffusion model, leading to CFP identifying the notion of concept signal. This kickstarted the project. ESL and HT defined the formal setup (Sec. 3). ESL hypothesized the point of sudden turn marks the emergence of hidden abilities, hence defining the project narrative around competence versus performance. AL, CFP, and ESL led the hidden emergence section (Sec. 4.4), with AL contributing the latent and MLP intervention and CFP contributing the input space one. MO and HT collaborated on the landscape theory (Sec. 4.3), with inputs from CFP. MO and ESL collaborated on the underspecification picture of learning dynamics, with inputs from HT (Sec. 5). Paper writing and literature review were led by ESL, with several inputs from all authors. Visualizations were led by MO and HT, with several inputs from CFP. Experiments on CelebA learning dynamics were led by CFP and MO, and corresponding latent intervention were led by AL.
Appendix A Schematic for Concept Space
Fig. 12 illustrates the concept space framework discussed in Sec. 3.
Figure 12: The concept space framework. $\mathcal{G}$ is an invertible data-generating that maps process maps sampled vectors $z\sim P(\mathcal{Z})$ (where $\mathcal{Z}\subset\mathbb{R}^{d}$) to an observation $x\in\mathbb{R}^{n}$ (an image in this work). The concept space, $\mathcal{S}:=\{z|z\sim P(\mathcal{Z})\}$, is defined as a space of all possible concept vectors generated from a compositional prior $P(\mathcal{Z})$. The mixing function $\mathcal{M}$ masks some concept variables, the masked concept variables are thus underspecified.
Appendix B Experimental Details
B.1 Synthetic data